# Chern insulators and topological flat bands in cavity-embedded kagome systems

### v1

@@ I INTRODUCTION
The manipulation of quantum phases via light-matter coupling has attracted significant attention in condensed matter physics [1, 2, 3, 4, 5, 6, 7, 8]. Optical cavities provide a powerful platform for tuning light-matter coupling between electrons and photons. Recent experiments have achieved strong light-matter interactions in cavities where the coupling strength exceeds a fraction of the bare cavity frequency, which is known as the ultrastrong couplling (USC) regime [9, 10, 11, 12, 13]. Because electronic states can be drastically modified in the USC regime, cavities provide a route for studying unconventional physics. Along with such experimental depelopments, recent theoretical studies have investigated various many-body phenomena in cavity quantum electrodynamics (QED), such as ferroelectricity [14, 15], superconductivity [16, 17, 18], quantum spin liquids [19, 20, 21], and Kondo effect [22, 23].
Furthermore, control of topological phases with the cavity has been explored extensively [24, 25, 26, 27, 11, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 13, 38]. For example, graphene has been predicted to host a Chern insulating phase in a circularly polarized cavity because Dirac points are gapped out by the light-matter interaction  [28, 36, 24, 25, 26, 33]. These topological phase transitions do not necessarily require strong light-matter interaction if time-reversal symmetry is broken [39]. A topological phase that appears only in the USC regime has yet to be found in graphene. So far, topological phases in the USC regime have remained elusive.
@@
The rest of this paper is organized as follows. In Sec. II, we introduce a model for kagome and honeycomb systems embedded in a circularly polarized cavity. We analyze band structures and topological phases of the model over a broad range of coupling strengths of the light-matter interaciton in Sec. III. To study topological edge states in the Chern insulating phases, we construct a low-energy tight-binding model and demonstrate the presence of the edge states in Sec. IV. Finally, a summary is given in Sec.V.
@@ II MODEL
where $m$ and e-e are an effective mass and the charge of an electron, respectively, and $V(\bm{r})$ is an arbitrary periodic potential. Here, $\hat{\bm{A}}=A_{0}(\bm{\varepsilon}\hat{a}+\bm{\varepsilon}^{*}\hat{a}^{\dagger})$ is the quantized vector potential with the mode amplitude $A_{0}$, where $\hat{{a}}\ (\hat{{a}}^{\dagger})$ is the annihilation (creation) operator of a photon with a cavity mode $\omega_{\mathrm{c}}$ and $\bm{\varepsilon}=(1,-i)^{\mathrm{T}}/\sqrt{2}$ is the circular polarization vector that breaks time-reversal symmetry. The coupling strength of the light-matter interaction is quantified by $g=eA_{0}\sqrt{\omega_{\mathrm{c}}/(m\hbar)}$. If $g/\omega_{c}\gtrsim 0.1$, this parameter range is typically called the USC regime.
We study kagome and honeycomb systems under a circularly polarized cavity field. Such systems can be described by the following muffin-tin potential:
@@
where $\bm{R}$ is a lattice vector given by the primitive vectors $\bm{a}_{1}=(a,\ 0)$ and $\bm{a}_{2}=(-a/2,\ \sqrt{3}a/2)$ with lattice constant $a$, and $\bm{L}=(0,a/\sqrt{3})$. As shown in Figs. 1(a) and 1(b), the potential $V(\bm{r})$ consists of barriers with height $V_{0}>0$ and radius $\rho_{\mathrm{H}}$ located at the lattice sites $\bm{R}$, and additional barriers with height $\alpha V_{0}$ and radius $\rho_{\mathrm{T}}$ centered at $\bm{R}\pm\bm{L}$. Here, $\alpha$ is a real dimensionless parameter, and the potential yields a kagome lattice for $\alpha=1$, and a honeycomb lattice for $\alpha=0$. Without vector potential, the band structures of the system with $\alpha=1$ have been studied as a kagome model [53].
@@
To confirm whether the potential in Eq. (2) imitates typical kagome and honeycomb lattices, we compute bulk energy bands in the absence of the light-matter interaction. We set $m=90\hbar^{2}/(a^{2}V_{0}),\hbar\omega_{\mathrm{c}}=0.2V_{0},\rho_{\mathrm{H}}=0.3a$, and $\rho_{\mathrm{T}}=\rho_{\mathrm{H}}/3$, and use these parameters throughout this paper. If $V_{0}=0.2~\mathrm{eV}$ and $a=5.5$ Å  are chosen, the effective mass is given by $m=1.1\times 10^{2}m_{\mathrm{e}}$ with the bare electron mass $m_{\mathrm{e}}$. Figure 1 (c) shows the three lowest energy bands for $\alpha=1$. We can find Dirac points at the $\mathrm{K}$ and $\mathrm{K}^{\prime}$ points and a nearly flat band with band degeneracy at the $\Gamma$ point. This band structure is consistent with the previous study [53], and is similar to that of typical kagome systems [40, 41, 42, 43]. We also show bulk bands for $\alpha=0$ in Fig. 1 (d). In this case, the nearly flat band is shifted to higher energy. In contrast, the Dirac points remain at the $\mathrm{K}$ and $\mathrm{K}^{\prime}$ points in the two lowest bands, which can be found in honeycomb lattice models such as graphene [54]. Hence, we call the systems with $\alpha=1$ and $\alpha=0$ kagome and honeycomb systems, respectively.
@@ III Chern insulating phases and topological flat bands induced by the cavity field
$C_{n}=\frac{1}{2\pi}\int_{\mathrm{BZ}}d^{2}k\ i\big(\langle\partial_{k_{x}}$ $\Psi_{n\bm{k}}|\partial_{k_{y}}\Psi_{n\bm{k}}\rangle$
$-\langle\partial_{k_{y}}\Psi_{n\bm{k}}|\partial_{k_{x}}\Psi_{n\bm{k}}\rangle\big),$ (3)
where $|\Psi_{n\bm{k}}\rangle$ is an electron-photon eigenstate for the $n$-th band. We label bands in ascending order from the bottom unless a band touching occurs. We characterize insulators with $n_{\mathrm{occ}}$ occupied bands using the total Chern number $C$, which is defined as the sum of the Chern numbers $C=\sum_{n=1}^{n_{\mathrm{occ}}}C_{n}$. In this paper, we investigate the case where the low-energy band structures shown in Figs. 1(c) and  1(d) are separated from other high-energy bands in the presence of light-matter interaction. This condition is satisfied if the photon energy $\hbar\omega_{c}$ is sufficiently larger than the bandwidth. For example, choosing $\hbar\omega_{c}=0.2V_{0}$ fulfills this condition, and the corresponding frequency is given by $\omega_{\mathrm{c}}=2\pi\times 9.7$ THz for $V_{0}=0.2~\mathrm{eV}$. We consider photon-number subspaces of up to 5 for diagonalizing the Hamiltonian in Eq. (1).
@@ III.1 Kagome system case: $\alpha=1$
We analyze the band topology in the kagome system by varying the coupling strength of the light-matter interaction. We illustrate Chern numbers of the kagome systems in Fig. 2 (a). When the coupling strength is introduced, the band degeneracy at the K, $\mathrm{K}^{\prime}$ and $\Gamma$ points is lifted [Fig. 2 (b)]. Due to the mass gap, topologically nontrivial bands with $(C_{1},C_{2},C_{3})=(1,0,-1)$ appear. The light-matter interaction opens a band gap between the first and second bands. Therefore, when the Fermi energy is inside the band gap between the first and second bands, a Chern insulator with a total Chern number of $C=1$ is realized. In contrast, when the light-matter interaction $g/\omega_{c}$ is weak, the third band, i.e., the flat band, can be topologically nontrivial with $C_{3}=-1$ while an indirect band overlap survives between the second and third bands.
Figure 2: (a) Chern numbers of the kagome system with the circularly polarized cavity field. Each band is topologically characterized by the Chern numbers $(C_{1},C_{2},C_{3})$. (b) Low-energy band structure at $g/\omega_{\mathrm{c}}=0.2$, and the band gap due to the cavity field at the K and $\Gamma$ points, respectively. (c)-(f) Low-energy band structures at $g/\omega_{\mathrm{c}}=1.65,1.8,2.02$ and $2.2$, respectively. (c) and (e) show the topological phase transitions at $g/\omega_{\mathrm{c}}=1.65$ and $2.02$, where the energy gap closes between the second and third bands at the $\Gamma$ point, and the first and second bands at the $\mathrm{K}$ point, respectively.
Moreover, topological phase transitions between different Chern insulating phases occur in the USC regime. When $g/\omega_{\mathrm{c}}\simeq 1.65$, the energy gap closes at the $\Gamma$ point, as shown in Fig. 2 (c). This gap closing gives rise to a topological phase transition that results in a topological band structure characterized by $(C_{1},C_{2},C_{3})=(1,-2,1)$. Thus, this phase transition changes the sign of the Chern number of the topological flat band characterized by $C_{3}$. Moreover, as shown in Fig. 2 (d), after the phase transition the indirect band overlap between the second and third bands becomes smaller as the light-matter interaction increases. Once the band gap entirely opens between the second and third bands, the system is a Chern insulator with $C=C_{1}+C_{2}=-1$ if the Fermi energy lies within this gap.
Next, we assume that the Fermi energy lies in the band gap between the first and second bands. When $g/\omega_{\mathrm{c}}\simeq 2.02$, the first and second bands touch at the K and $\mathrm{K}^{\prime}$ points [Fig. 2 (e)], which leads to topological bands with $(C_{1},C_{2},C_{3})=(-1,0,1)$. Thus, the system exhibits a Chern insulating phase with $C=C_{1}=-1$. We show a band structure with $(C_{1},C_{2},C_{3})=(-1,0,1)$ in Fig. 2(f). As a result, the system undergoes topological phase transitions between the phases with $C=1$ and C=1C=-1 in the USC regime, when the Fermi energy lies in the lower band gap. Therefore, we find that the band topology of the kagome systems can be changed by varying the coupling strength. Even though we change parameters in the model, such phase transitions can occur as long as the band gap above the third band remains open.
@@ III.2 Honeycomb system case: $\alpha=0$
We next study band topology in the honeycomb system with the cavity field. Figure 3 shows its band structure. The light-matter coupling gives a mass gap to the Dirac points at the K and $\mathrm{K}^{\prime}$ points, which induces bands with $(C_{1},C_{2})=(1,-1)$. Thus, a Chern insulating phase with $C=1$ is realizable when only the lowest energy band is occupied. However, unlike the kagome system, no topological phase transition happens even though the light-matter coupling becomes stronger, which is consistent with previous works [28, 36]. This behavior does not change, as long as the band gap above the second band remains open.
@@ IV Topological edge states
Figure 3: Low-energy band structure in the honeycomb system at the coupling strength $g/\omega_{\mathrm{c}}=0.5$.
Figure 4: (a)-(c) Comparison of bulk band structures from the exact analysis and the tight-binding model at $g/\omega_{\mathrm{c}}=0.2,1.8$, and $2.9$, which are topologically characterized by $(C_{1},C_{2},C_{3})=(1,0,-1),(1,-2,1)$, and (1,0,1)(-1,0,1), respectively. We adjust the lowest energy eigenvalue obtained from the tight-binding model at the K point to match that obtained from the exact analysis because the energy shift does not change the band topology. (d)-(f) Topological edge states for $g/\omega_{\mathrm{c}}=0.2,1.8$, and $2.9$, respectively, with the periodic boundary condition in the $x$-direction and open boundary condition in the $y$-direction. Red (green) solid lines indicate gapless edge modes localized at the top (bottom) edge.
@@
Then, the Hamiltonain in Eq. (1) is transformed into
@@
where $m_{\mathrm{eff}}=m(1+g^{2}/\omega_{\mathrm{c}}^{2})$ is the effective mass dressed by cavity photons and $\Omega=\omega_{c}(1+g^{2}/\omega_{\mathrm{c}}^{2})$ is the renormalized cavity frequency, $\bm{e}_{z}$ is the unit vector in the $z$-direction. In this AD frame, the effect of the quantum electromagnetic field is incorporated into the electronic Hamiltonian through the effective mass and the modified periodic potential. Since we focus on low-energy bands, we project the AD-frame Hamiltonian onto the zero-photon subspace. As a result, we obtain the effective low-energy Hamiltonian given by

### v2

@@ I INTRODUCTION
The manipulation of quantum phases via light-matter coupling has attracted significant attention in condensed matter physics [1, 2, 3, 4, 5, 6, 7, 8]. Optical cavities provide a powerful platform for tuning light-matter coupling between electrons and photons. Recent experiments have achieved strong light-matter interactions in cavities where the coupling strength exceeds a fraction of the bare cavity frequency, which is known as the ultrastrong coupling (USC) regime [9, 10, 11, 12, 13, 14]. Because electronic states can be drastically modified in the USC regime, cavities provide a route for studying unconventional physics. Along with such experimental developments, recent theoretical studies have investigated various many-body phenomena in cavity quantum electrodynamics (QED), such as ferroelectricity [15, 16], superconductivity [17, 18, 19], quantum spin liquids [20, 21, 22], and Kondo effect [23, 24].
Furthermore, control of topological phases with the cavity has been explored extensively [25, 26, 27, 28, 12, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 14, 42, 43, 44]. For example, graphene has been predicted to host a Chern insulating phase under a circularly polarized cavity field because Dirac points are gapped out by the light-matter interaction [29, 39, 25, 26, 27, 36]. These topological phase transitions do not necessarily require strong light-matter interaction if time-reversal symmetry is broken [45]. A topological phase that appears only in the USC regime has yet to be found in graphene. So far, topological phases in the USC regime have remained elusive.
@@
The rest of this paper is organized as follows. In Sec. II, we introduce a model for kagome and honeycomb systems embedded in a chiral cavity. We analyze band structures and topological phases of the model over a broad range of coupling strengths of the light-matter interaction in Sec. III. To study topological edge states in the Chern insulating phases, we construct a low-energy tight-binding model and demonstrate the presence of the edge states in Sec. IV. Finally, a summary is given in Sec. V.
@@ II MODEL
where $m$ and e-e are an effective mass and the charge of an electron, respectively, and $V(\bm{r})$ is an arbitrary periodic potential. In this paper, we neglect both direct and cavity-mediated electron-electron interactions, and calculate band structures  [^1] To calculate band structures in momentum space, we apply the transformation $\hat{H}_{\mathrm{C}}(\bm{k})=e^{-i\bm{k}\cdot\bm{r}}\hat{H}_{\mathrm{C}}e^{i\bm{k}\cdot\bm{r}}$. We then diagonalize $\hat{H}_{\mathrm{C}}(\bm{k})$ by expanding the eigenstates in the basis of the product state of a plane wave state and a photon number eigenstate $|\bm{\mathcal{K}}\otimes n\rangle$, where $\bm{\mathcal{K}}$ is a reciprocal lattice vector and $n$ is the photon number, and obtain the corresponding energy eigenvalues . Here, $\hat{\bm{A}}=A_{0}(\bm{\varepsilon}\hat{a}+\bm{\varepsilon}^{*}\hat{a}^{\dagger})$ is the quantized vector potential with the mode amplitude $A_{0}$, where $\hat{{a}}\ (\hat{{a}}^{\dagger})$ is the annihilation (creation) operator of a photon with a cavity mode $\omega_{\mathrm{c}}$ and $\bm{\varepsilon}=(1,-i)^{\mathrm{T}}/\sqrt{2}$ is the circular polarization vector that breaks time-reversal symmetry. Such a circularly polarized cavity mode can be realized in chiral cavities [60]. The coupling strength of the light-matter interaction is quantified by $g=eA_{0}\sqrt{\omega_{\mathrm{c}}/(m\hbar)}$ [61, 29]. We vary the coupling strength by changing the vector potential. If $g/\omega_{\mathrm{c}}\gtrsim 0.1$, this parameter range is typically called the USC regime.
We study kagome and honeycomb systems coupled to a circularly polarized mode in a chiral cavity, as illustrated in Fig. 1. Such systems can be described by the following muffin-tin potential:
@@
where $\bm{R}$ is a lattice vector given by the primitive vectors $\bm{a}_{1}=(a,\ 0)$ and $\bm{a}_{2}=(-a/2,\ \sqrt{3}a/2)$ with lattice constant $a$, and $\bm{L}=(0,a/\sqrt{3})$. As shown in Figs. 2(a) and 2(b), the potential $V(\bm{r})$ consists of barriers with height $V_{0}>0$ and radius $\rho_{\mathrm{H}}$ located at the lattice sites $\bm{R}$, and additional barriers with height $\alpha V_{0}$ and radius $\rho_{\mathrm{T}}$ centered at $\bm{R}\pm\bm{L}$. Here, $\alpha$ is a real dimensionless parameter, and the potential yields a kagome lattice for $\alpha=1$, and a honeycomb lattice for $\alpha=0$. Without vector potential, the band structures of the system with $\alpha=1$ have been studied as a kagome model [62]. We note that tunable honeycomb and kagome band structures have been experimentally realized in GaAs-based two-dimensional electron gases (2DEGs) using artificially patterned potentials [63, 64, 65].
Figure 1: Schematic drawing of a kagome material embedded in a chiral cavity. The arrow represents a circularly polarized cavity mode.
@@
To confirm whether the potential in Eq. (2) imitates typical kagome and honeycomb lattices, we compute bulk energy bands in the absence of the light-matter interaction. We set $m=90\hbar^{2}/(a^{2}V_{0}),\hbar\omega_{\mathrm{c}}=0.2V_{0},\rho_{\mathrm{H}}=0.3a$, and $\rho_{\mathrm{T}}=\rho_{\mathrm{H}}/3$, and use these parameters throughout this paper. For example, if $V_{0}=10.2~\mathrm{meV}$ and $a=100~\mathrm{nm}$ are chosen, we obtain $m=0.067m_{\mathrm{e}}$ with the bare electron mass $m_{\mathrm{e}}$, which is in line with the effective mass of GaAs-based 2DEGs with patterned potential [65]. Figure 2 (c) shows the three lowest energy bands for $\alpha=1$. We can find Dirac points at the $\mathrm{K}=2\pi/a(1/3,1/\sqrt{3})$ and $\mathrm{K}^{\prime}=2\pi/a(2/3,0)$ points and a nearly flat band with band degeneracy at the $\Gamma=(0,0)$ point. This band structure is consistent with the previous study [62], and is similar to that of typical kagome systems [46, 47, 48, 49]. We also show bulk bands for $\alpha=0$ in Fig. 2 (d). In this case, the nearly flat band is shifted to higher energy. In contrast, the Dirac points remain at the $\mathrm{K}$ and $\mathrm{K}^{\prime}$ points in the two lowest bands, which can be found in honeycomb lattice models such as graphene [66]. Hence, we call the systems with $\alpha=1$ and $\alpha=0$ kagome and honeycomb systems, respectively.
@@ III Chern insulating phases and topological flat bands induced by the cavity field
Figure 3: (a) Chern numbers of the kagome system with the circularly polarized cavity field. Each band is topologically characterized by the Chern numbers $(C_{1},C_{2},C_{3})$. (b) Low-energy band structure at $g/\omega_{\mathrm{c}}=0.2$, and the band gap due to the cavity field at the $\mathrm{K}$ and $\Gamma$ points, respectively. (c)-(f) Low-energy band structures at $g/\omega_{\mathrm{c}}=1.65,1.8,2.02$ and $2.2$, respectively. (c) and (e) show the topological phase transitions at $g/\omega_{\mathrm{c}}=1.65$ and $2.02$, where the energy gap closes between the second and third bands at the $\Gamma$ point, and the first and second bands at the $\mathrm{K}$ point, respectively. The color of the bands indicates the expectation value of the photon number in panels (b)-(f).
@@
$C_{l}=\frac{1}{2\pi}\int_{\mathrm{BZ}}d^{2}k\ i\big(\langle\partial_{k_{x}}$ $\Psi_{l\bm{k}}|\partial_{k_{y}}\Psi_{l\bm{k}}\rangle$
$-\langle\partial_{k_{y}}\Psi_{l\bm{k}}|\partial_{k_{x}}\Psi_{l\bm{k}}\rangle\big),$ (3)
where $|\Psi_{l\bm{k}}\rangle$ is an electron-photon eigenstate for the $l$-th band. We label bands in ascending order from the bottom unless a band touching occurs. We characterize insulators with $l_{\mathrm{occ}}$ occupied bands using the total Chern number $C$, which is defined as the sum of the Chern numbers $C=\sum_{l=1}^{l_{\mathrm{occ}}}C_{l}$. In this paper, we investigate the case where the low-energy band structures shown in Figs. 2(c) and  2(d) are separated from other high-energy bands in the presence of light-matter interaction. This condition is satisfied if the photon energy $\hbar\omega_{\mathrm{c}}$ is sufficiently larger than the bandwidth. For example, choosing $\hbar\omega_{\mathrm{c}}=0.2V_{0}$ fulfills this condition, and the corresponding frequency is given by $\omega_{\mathrm{c}}=2\pi\times 494$ GHz for $V_{0}=10.2~\mathrm{meV}$.
In our parameter regime, we also compute the expectation value of the photon number $\langle\hat{a}^{\dagger}\hat{a}\rangle=\braket{\Psi_{l\bm{k}}|\hat{a}^{\dagger}\hat{a}|\Psi_{l\bm{k}}}$ for the low-energy bands (See Figs. 3 and 4). The expectation value remains well below 1 for $g/\omega_{\mathrm{c}}\leq 3$. Thus, we consider photon-number subspaces of up to 5 for diagonalizing the Hamiltonian in Eq. (1), which is sufficient to investigate the band topology.
@@ III.1 Kagome system case: $\alpha=1$
We analyze the band topology in the kagome system by varying the coupling strength of the light-matter interaction. We illustrate Chern numbers of the kagome systems in Fig. 3 (a). When the coupling strength is introduced, the band degeneracy at the $\mathrm{K}$, $\mathrm{K}^{\prime}$ and $\Gamma$ points is lifted [Fig. 3 (b)]. Due to the mass gap, topologically nontrivial bands with $(C_{1},C_{2},C_{3})=(1,0,-1)$ appear.
The light-matter interaction opens a band gap between the first and second bands. Therefore, when the Fermi energy is inside the band gap between the first and second bands, a Chern insulator with a total Chern number of $C=1$ is realized. In contrast, when the light-matter interaction $g/\omega_{\mathrm{c}}$ is weak, the third band, i.e., the flat band, can be topologically nontrivial with $C_{3}=-1$ while an indirect band overlap survives between the second and third bands.
Moreover, topological phase transitions between different Chern insulating phases occur in the USC regime. When $g/\omega_{\mathrm{c}}\simeq 1.65$, the energy gap closes at the $\Gamma$ point, as shown in Fig. 3 (c). This gap closing gives rise to a topological phase transition that results in a topological band structure characterized by $(C_{1},C_{2},C_{3})=(1,-2,1)$. Thus, this phase transition changes the sign of the Chern number of the topological flat band characterized by $C_{3}$. Moreover, as shown in Fig. 3 (d), after the phase transition, the indirect band overlap between the second and third bands becomes smaller as the light-matter interaction increases. Once the band gap entirely opens between the second and third bands, the system is a Chern insulator with $C=C_{1}+C_{2}=-1$ if the Fermi energy lies within this gap.
Next, we assume that the Fermi energy lies in the band gap between the first and second bands. When $g/\omega_{\mathrm{c}}\simeq 2.02$, the first and second bands touch at the $\mathrm{K}$ and $\mathrm{K}^{\prime}$ points [Fig. 3 (e)], which leads to topological bands with $(C_{1},C_{2},C_{3})=(-1,0,1)$. Thus, the system exhibits a Chern insulating phase with $C=C_{1}=-1$. We show a band structure with $(C_{1},C_{2},C_{3})=(-1,0,1)$ in Fig. 3(f). As a result, the system undergoes topological phase transitions between the phases with $C=1$ and C=1C=-1 in the USC regime, when the Fermi energy lies in the lower band gap. Therefore, we find that the band topology of the kagome systems can be changed by varying the coupling strength. Even though we change parameters in the model, such phase transitions can occur as long as the band gap above the third band remains open.
@@ III.2 Honeycomb system case: $\alpha=0$
We next study band topology in the honeycomb system with the cavity field. Figure 4 shows its band structure. The light-matter coupling gives a mass gap to the Dirac points at the $\mathrm{K}$ and $\mathrm{K}^{\prime}$ points, which induces bands with $(C_{1},C_{2})=(1,-1)$. Thus, a Chern insulating phase with $C=1$ is realizable when only the lowest energy band is occupied. However, unlike the kagome system, no topological phase transition happens even though the light-matter coupling becomes stronger, which is consistent with previous works [29, 39]. This behavior does not change, as long as the band gap above the second band remains open.
Here we compare the band topology of the cavity-embedded kagome and honeycomb systems in our parameter regime. In the kagome system, the three-band structure enables gap closings between different pairs of bands as the coupling strength is increased. Such gap closings allow Chern numbers $C_{l}$ to be redistributed among the three bands. By contrast, in the honeycomb model, the gap closing occurs only between the two bands. In view of the number of the bands, this behavior is more likely in the kagome system than in the minimal two-band honeycomb model. As a result, the cavity-embedded kagome system supports various topological bands.
@@ IV Topological edge states
Figure 4: Low-energy band structure in the honeycomb system at the coupling strength $g/\omega_{\mathrm{c}}=0.5$. The color of the bands indicates the expectation value of the photon number.
Figure 5: (a)-(c) Comparison of bulk band structures from the exact analysis and the tight-binding model at $g/\omega_{\mathrm{c}}=0.2,1.8$, and $2.9$, which are topologically characterized by $(C_{1},C_{2},C_{3})=(1,0,-1),(1,-2,1)$, and (1,0,1)(-1,0,1), respectively. We adjust the lowest energy eigenvalue obtained from the tight-binding model at the $\mathrm{K}$ point to match that obtained from the exact analysis because the energy shift does not change the band topology. (d)-(f) Topological edge states for $g/\omega_{\mathrm{c}}=0.2,1.8$, and $2.9$, respectively, with the periodic boundary condition in the $x$-direction and open boundary condition in the $y$-direction. Red and green solid lines represent gapless edge modes, and the color difference indicates that these modes are localized at opposite edges.
@@
Then, the Hamiltonian in Eq. (1) is transformed into
@@
where $m_{\mathrm{eff}}=m(1+g^{2}/\omega_{\mathrm{c}}^{2})$ is the effective mass dressed by cavity photons and $\Omega=\omega_{\mathrm{c}}(1+g^{2}/\omega_{\mathrm{c}}^{2})$ is the renormalized cavity frequency, $\bm{e}_{z}$ is the unit vector in the $z$-direction. In this AD frame, the effect of the quantum electromagnetic field is incorporated into the electronic Hamiltonian through the effective mass and the modified periodic potential. Since we focus on low-energy bands, we project the AD-frame Hamiltonian onto the zero-photon subspace. As a result, we obtain the effective low-energy Hamiltonian given by